### v1

# Human-interpretable clustering of short-text using large language models

Abstract
Large language models have seen extraordinary growth in popularity due to their human-like content generation capabilities. We show that these models can also be used to successfully cluster human-generated content, with success defined through the measures of distinctiveness and interpretability. This success is validated by both human reviewers and ChatGPT, providing an automated means to close the ‘validation gap’ that has challenged short-text clustering. Comparing the machine and human approaches we identify the biases inherent in each, and question the reliance on human-coding as the ‘gold standard’. We apply our methodology to Twitter bios and find characteristic ways humans describe themselves, agreeing well with prior specialist work, but with interesting differences characteristic of the medium used to express identity.
1 Introduction
Short text is playing an increasingly important role in human expression and interaction, due to the widespread use of social media platforms and messaging services such as X (formerly Twitter), Weibo, WhatsApp, Instagram and Reddit. The enormous quantities of data produced by users of these platforms holds the promise of not just real-time identification of events [1] and current opinions [2], but also a deeper understanding of the drivers of information flow between the users [3]. A first step in engaging with the large data sets is to typically reduce the complexity, by clustering the text data into similar groups [4]. However, short text clustering is challenging, due to the limited contextual information available in a single piece of text, and the low incidence of word co-occurrence between pieces of text [5, 6].
The possible applications of success has led to a focus in the machine learning community on clustering, with an increasing number of methods developed to provide a deeper understanding of large collections of short text data [7, 8]. However, a major criticism facing current clustering approaches is that the resulting clusters appear largely uninterpretable by humans [9]. In addition, automated metrics used to quantify clustering success appear to be poorly correlated with interpretability [10]. There is thus a two-fold challenge, the development of an approach which can lead to interpretable clusters of short text data, and an automated approach to stand as a proxy for humans interpreting the resulting clusters.
Approaches to the clustering of short text can be placed under three broad categories: frequency-based approaches, embedding methods, and the use of deep learning techniques [7]. Traditionally, approaches to text involve looking for topics which span the range of documents, in a process known as topic modelling [8]. Typically this involves building a probabilistic model based on observed word frequencies [11], and then for short text, assuming that each document belongs to only one topic [8]. Documents in a given topic are then taken to form a cluster [12]. The most commonly employed topic modelling approach is latent Dirichlet allocation [11], due to its ease of use, however the severe sparsity of short text data is known to lead to poor performance [6].
Embedding approaches instead seek to represent the text data in a vector space [7], where standard clustering techniques such as k-means or Gaussian mixture modelling can be applied [13]. We consider in this class methods such as doc2vec [14], which establish an embedding space by training on the data set of interest. This is in contrast to embeddings created using deep learning methods, such as the Bidirectional Encoder Representations in Transformers (BERT) [15, 16]. In a transformer-based approach, the model is built by training on a large external data set, and then the data set of interest is embedded in the vector space created using the training set.
While transformers have shown early promise for text clustering [16], they have seen spectacular recent success in text generation, thanks to large language models such as ChatGPT [17]. ChatGPT has been rapidly embraced by the public, but large language models more broadly have also seen applications in areas as diverse as finance [18], health [19], law [20] and academia [21]. The transformer-based architectures underpinning these models typically have billions of parameters and are trained on hundreds of GB of text data [22]. While the focus of these models has been on producing human-like content [23], more quantitative applications are starting to emerge, such as sentiment analysis [24] and text annotation [25]. The extraordinary capabilities of these models to produce extended examples of human-like text raises the prospect that they might also be used to group together similar pieces of text, and potentially even be used to interpret the resulting groups.
In this work we use large language models (LLMs) to cluster short text and to interpret the resulting clusters. Out of the many possible clustering techniques available, we compare the LLM with LDA, as a prominent example of topic modelling, and doc2vec as a non-deep-learning approach. Despite the limitations of LDA the method is still widely used [26] and was recently employed in a comparative work on short text clustering methods [12]. Doc2vec was used in the same comparative work, and found to provide the best results when judged using a set of ‘gold standard’ labels of short text documents [12]. In contrast, we consider the case of unlabelled data, and seek to quantify clustering success through human interpretability.
A complication which arises when framing success in cluster creation using human interpretability, is that humans have biases and limitations when faced with the tasks of clustering (also called categorization, the identification of unlabelled groups of similar objects) and classification (labelling of already existing clusters) [27]. Humans appear to cluster preferentially along only one selected dimension in high dimensional data [28], and the dimension chosen depends on the experiences and biases of an individual performing the clustering [29]. Human-labelled data is typically taken as a base truth, or gold standard, however these limitations suggest that, at least for cluster labels, there may be no such thing as a gold standard. One of the goals of this work is to determine a measure of clustering success in the presence of this ambiguity.
We use as a test case for human-interpretable automated short-text clustering a domain familiar from lived experience, that of human identity. The nature of how humans self-identify is complex, however, twelve categories have been found in analysis of identity-related nouns in the English language: occupation, political, kinship, religion, biosocial, ethnicity, corporal, leisure, sexuality, stigma, esteemed, and other [30]. When seeking to classify identity based on short text, we expect there to be substantial ambiguity, making classification challenging for automated approaches and for human reviewers. However, success in this domain promises insights into social media users. Knowledge of identity can be predictive of a person’s behaviour in many areas of life. For instance, political identity can predict a person’s beliefs and actions, such as believing in conspiracy theories or participating in activism for climate change[31, 32]. Humans are interacting to an ever larger degree via social media, and these interactions are typically mediated by short text. With our methodology we show that it is possible to infer information about the people producing the text, when performed at the large data scale.
We focus specifically on a short text data set consisting of ‘Twitter bios’, which contains user responses to the general prompt ‘Describe yourself’. Users have 160 characters to respond to this prompt, resulting in highly variable short text. We begin by using the three methods of LDA, doc2vec and an LLM to find clusters in this data set. We then ask human reviewers to interpret and rate the resulting clusters. We examine what constitutes a ‘good’ cluster in light of the cluster characteristics and the human reviews and look for possible metrics that can provide a quantitative measure of success. We then turn to the possibility of using ChatGPT to act as a proxy for the human reviewers. We examine the results of both the machine and human approaches in light of what is known about human identity and conclude with a proposal of a method for finding and interpreting short text clusters.
2 Results
We use a short-text dataset consisting of 38,639 Twitter user bios from users who used the keywords ‘trump’ or ‘realDonaldTrump’ on September 3rd and 4th 2020 (see Methods for further details). This limits the domain of users to those with some interest in US politics. In Fig. 1 we get an overview of the nature of the bios by examining the top fifty word co-occurrences in the bios. In this ‘top-word co-occurrence graph’ the sizes of the nodes represent the frequency of word occurrence, and the edges represent the number of pairwise co-occurrences. The US-centric nature of the dataset is immediately evident with the large US flag node. The political nature of the users is also clear with both left-leaning (e.g. resist, blue wave emoji, liberal) and right-leaning words (e.g. maga, wwg1wga, trump2020) sitting in clusters. While the bios appear to be dominated by partisan political members, there is also evidence of other identity flags, with a family-oriented cluster of co-occurrences and an animal-lover cluster, so we see at this coarse level already the possibility of multiple ways of users self-identifying.
Figure 1: A top-word co-occurrence graph, illustrating the co-occurrence relationships among the top 50 words in the dataset. Each node, sized according to its relative frequency, is connected only to the two words with which it most frequently co-occurs. The edge thickness represents the frequency of these co-occurrence relationships. Notably, the ’US’ node, symbolized by the American flag emoji.
Figure 2: A cluster overlap graph in which nodes represent the clusters by model (Top row: LLM; Middle row: LDA; Bottom row: doc2vec). Lines represent overlap between cluster pairs, as measured by the number of bios in common in a pair. A threshold of 1190 overlapping bios is set for the appearance of a line, which represents the top 5% of cluster overlaps. Thicker lines represent greater numbers of bios in common. Maximum overlap is 2,638 between the ‘Right’ cluster of the LLM and of LDA. Cluster names are given by the authors (see text).
Given the evidence of partisan political users in this set, we would expect any clustering of the bios should contain clusters separated along ideological lines. We proceed with the three methods identified earlier: LDA, doc2vec and an LLM (for implementation details, see the Methods section).
In each case we restrict the number of clusters to 10, so that the methods can be more equally compared. With ten clusters we aim to find a balance between complexity (number of clusters to identify) and distinctiveness (how well-defined each cluster is), such that naming is possible for reviewers. An interesting further direction is to consider how the results vary with the number of clusters.
Fig. 2 graphically represents the partitions by row for the three methods (top: LLM; middle: LDA; bottom: doc2vec). The size of a node reflects the cluster size, the thickness of lines connecting nodes indicates the number of bios in common between the clusters, while the labels have been given by the researchers (see Methods). Only the LLM clusters could be comprehensively named. It is evident in Fig. 2 that there is significant variability in the sizes of the clusters produced, with the doc2vec method particularly varied. Somewhat surprisingly, given the political nature of the dataset, only one cluster appears to show any appreciable overlap between the methods, as indicated by the thicker line linking a cluster in the LDA and LLM methods. This appears to be for politically right-leaning users. Overall, the methods appear to have created largely distinct partitions of the short-text data.
A key measure of interpretability is the ability of humans to easily name a cluster [9] and see agreement between top words of a cluster and sample documents. Coherence is a commonly-used automated metric, however its performance as a useful measure has come under recent scrutiny [33]. Mindful of earlier guidelines on human interpretability of clusters [9] we recruited 39 participants to rate: the coherence of a set of 20 sample bios chosen randomly from a cluster, the coherence of the top 10 words from each cluster, and the coherence between the sample bios and the top words (see Methods for details). We also asked the reviewers to name each cluster (if possible) and rate their confidence that the chosen name correctly describes the cluster. As a null model we included also randomly generated clusters, with top words and sample bios from these random clusters provided to the reviewers. The results of the model comparison can be seen in Fig. 3.
Figure 3: Ordinal Regression Analysis of Clusters created by LDA, Doc2vec, and LLM illustrating the outcomes of an ordinal regression model applied to the ratings of 39 reviewers. The reviewers assessed Twitter bio clusters and were asked to rate the coherence of the clusters across four categories: (a) Confidence in Cluster Name, (b) Coherence of Top Words, (c) Coherence of Sample Bios, and (d) Coherence between the top words and sample bios. Each panel (a-d) represents the density distribution ($\hat{Y}$ ) of ratings in each category, showcasing the statistical modeling of ordered categorical data. The reviewer ratings of the random clusters are set to 0 so any value greater than 0 is performing better than random. The vertical lines identify the transitions between values on the likert scales used by the reviewers. We see that the LLM has consistently been scored at 4 on all measures.
As the human reviewers have rated their responses on a likert scale, we use ordinal regression [34] with respect to the baseline found for the random set, which is set to 0 (see Methods for further details). We see that the LLM has performed consistently better than the other two methods on all metrics. LDA performed consistently better than random, while doc2vec performed consistently worse. Inspection of the doc2vec clusters reveals that the clustering approach has indeed made use of features that are largely invisible to human reviewers, such as a cluster of bios using low-frequency words. Interestingly doc2vec was found to be successful at clustering when a gold standard is present [12], but in this case there is no gold standard. The results are consistent with the finding that traditional clustering approaches produce clusters that are often uninterpretable by humans [9].
Figure 4: The median reviewer score for each cluster across the four categories: (a) Confidence in Cluster Name, (b) Coherence of Top Words, (c) Coherence of Sample Bios, and (d) Coherence between the top words and sample bios. Error bars represent the first and third quartile scores. All clusters with the U designation were unable to be named by the authors of this paper.
When we look at the performance by cluster in Fig. 4, we find the results are more nuanced. In Fig. 4(a) we see that while the reviewers were usually confident in naming clusters produced by the LLM, some of the clusters were less clear, and one cluster could not be named at all. The cluster names provided on the $x$-axis have been created by the authors based on observation of the reviewer-provided names and a deeper review of the clusters (consistent with the names used in Fig. 2, see Methods for further details). The ‘General’ cluster could not be named, and the ‘Twitter’, ‘Quotes’ and ‘Emojis’ clusters were also challenging for the reviewers. We examine the names provided by reviewers further below. Interestingly, as seen in Fig. 4(b), coherence of top words, as rated by the reviewers, is not a good predictor of confidence in naming, with little distinction detected between the models, and even the randomly generated clusters rated highly. Looking at Fig. 4(c,d), the LLM appears to have aggregated bios such that they appear more coherent to the reviewers, as well as producing clusters with greater coherence between the bios and the top words. This suggests that top words can play a role in naming, as is common practice with topic modelling [9], however the high ratings also assigned to random clusters suggests that human estimations of coherence do not provide a measure of clustering success.
While interpretability has been identified as important, we suggest that clusters should also be distinctive. A subset of a randomly generated cluster may appear coherent, and even interpretable, to a human reviewer, due to the well-known human bias of seeing patterns in noise [35]. However, when dealing with large randomly-generated clusters, any fluctuations are expected to average out and the clusters will appear similar to each other. We need a measure to capture this degree of distinctiveness between clusters. To do so, we make use of the method of keyword analysis from corpus linguistics. We identify the number of keywords special to a cluster (see Methods) and plot this versus human confidence in naming, see Fig. 5. We find that all the clusters produced by the LLM have large numbers of distinctive keywords, even the clusters which human reviewers were unable to interpret (such as the ‘General’ cluster). In contrast, all the randomly generated clusters have zero, or close to zero, distinctive keywords, as expected for a complete mixing of the data set. The methods of LDA and doc2vec lead to some clusters with distinctive keywords, but much less distinctiveness than the LLM approach. We find therefore that the large language model has allowed for the creation of not just more human interpretable clusters, but also clusters which are more distinctive relative to each other.
Figure 5: The number of keywords in each cluster, where keywords are determined by comparing the expected frequency of words in a cluster against their actual frequency, using a Bayesian factor greater than 10 to assess if the difference is statistically significant. The names of each cluster are the names given by the authors and the prefix U indicates that a cluster was not able to be named. Colors identify the clustering model used and are consistent with the color code used in Fig. 4. The LLM clusters have significantly more keywords than the clusters found using the other methods.
Ideally, we would have an automated approach in which clusters are created and then a metric is used to quantify clustering success. We have identified an automated approach for measuring distinctiveness, we would also like to have a metric which correlates with human interpretability. The problems facing the use of automated metrics have been widely identified [10, 36, 9, 33, 8], with the popular metric of coherence singled out [9, 33]. While there are numerous methods used to calculate coherence [9] we choose two of the most common: CV coherence and UMASS coherence [37, 38, 33]. We also take advantage of the LLM’s ability to convert text into a vector space and use cosine similarity, Euclidean distance and the silhouette score to provide a measure of clustering success. In Fig. 6 we examine the correlation between these metrics and the ratings provided by the reviewers. We see that the correlation is generally poor. The Mean Standard Deviation metric, which serves as a proxy for cluster width in the embedded space, shows the best correlation with the human ratings, followed by the silhouette score. The coherence measures show little correlation with the human measures, including the human ratings of coherence. We find that these coherence measures also show poor correlation with each other (see Fig. S1), so failing a posited criterion of interpretability, that different coherence measures should correlate [9]. Our results are therefore consistent with findings elsewhere that, despite being widely used, the metric of coherence is not useful when measuring human interpretability [33]. We note that the number of keywords also does not correlate with the human measures, however as discussed earlier, we argue this metric provides a measure orthogonal to interpretability, and a lack of correlation is consistent with this.
Figure 6: Boxplots showing the Spearman Rank Correlation between the reviewer provided ratings and the six automated metrics of mean cluster standard deviation, silhouette score, Euclidean distance, number of keywords, CV coherence and UMass coherence (x-axis labels), with reviewer ratings on (a) Confidence in naming; (b) Coherence of top words; (c) Coherence of sample bios and (d) Coherence between top words and sample bios. Coherence and keywords correlate poorly with reviewer ratings. Mean standard deviation appears to provide the best correlation with the ratings. The large variability in correlation across reviewers is evident, with outliers identified with solid circles, i.e., some reviewers correlated well with some measures, while others showed no correlation or negative correlation for some measures.
So far we have focused on the quantification of the clusters provided by the reviewers and the automated metrics, however typically it is nature of the classification which is most interesting when examining clusters. The consistency of names provided by reviewers is also a measure of interpretability [9]. To this end, we turn now to the names of the clusters. The names given to some of the clusters in Fig. 4 are given by the authors based on top words and an extended set of bios in each cluster, as well as the names given by the reviewers. However more interesting is to see the consistency in naming provided by the reviewers. In Fig. 7(a) we show the most popular words appearing by cluster in the reviewer LLM cluster names. The x-axis tick labels are again the researcher-given names for the associated clusters. We see that the Right Leaning cluster has been consistently named by reviewers to be a ‘trump’ related cluster, and the Relational cluster is similarly annotated by the word ‘family’. This is consistent with the coarse clustering approach shown in Fig. 1, where both right-leaning and relational clusters were evident in the top words. The existence of ‘Relational’, ‘Occupation’ and ‘Political’ self-identities are consistent with the broader study of human identity [30]. Other clusters are less obvious and are perhaps specific to the medium of the social network. A sizeable proportion of Twitter account bios reference the use of Twitter as a primary activity, and were picked up and named by the reviewers. Others however appear to have been challenging for the reviewers to name, such as the ‘General’, ‘Quotes’ and ‘Emojis’ clusters. We seek now to identify whether these were difficult to name because they are poorly-defined clusters, or because the classification task was difficult due to the limited information available to the reviewers.
Figure 7: The top five words by fraction of appearance used by (a) reviewers and (b) ChatGPT to name the clusters created by the LLM. Along the x-axis are the names given to each cluster by the authors of this paper. We see that the words used to describe the clusters are largely consistent between ChatGPT and the reviewers, however there are cluster-dependent distinctions revealing human and machine limitations as discussed in the text.
To explore this problem, we contrast the reviewer names, Fig. 7(a), with those produced by ChatGPT-4.0, Fig. 7(b). We find that ChatGPT provides names for all the clusters, unlike many of the human reviewers who selected ‘None’ for some of the clusters. The names ChatGPT has given look consistent with the names provided by the reviewers, and those used by the researchers. ‘General’ and ‘Quotes’ have both been given appropriate names. This suggests that the human reviewers struggled to identify the underlying threads in the clusters of bios. This is consistent with a bias identified in human classification, in which humans tend to seek a low dimensional representation for data when they are asked to perform a classification task (e.g. choosing to focus only on expressions of political ideology) [29]. ‘Quotes’ do not easily fit into this representation, although we note that some reviewers did detect this feature in the set of bios. The ‘General’ class has been picked up by ChatGPT, and such a class is known to exist in expressions of human identity [30], but this class was overwhelmingly left unnamed by the reviewers. Interestingly, the ‘Emojis’ cluster was named by ChatGPT according to its content rather than the medium. This is perhaps unsurprising, given the tool focuses on word usage, and so has provided names based on the content rather than the form of the content. We conclude therefore that clusters that weren’t named by the reviewers, could be named by ChatGPT, suggesting human bias rather than limited information may be involved, however we also found that when faced with finding a ‘meta’ name for a cluster, the humans performed better than ChatGPT.
We are now ready to implement an automated approach for quantifying clustering success. In particular we wish to quantify the performance of ChatGPT with respect to the human reviewers. We simulate this by asking ChatGPT to name clusters multiple times and using the resulting names of both the human reviewers and ChatGPT to calculate consistency (a measure of interpretability) and distinctiveness of the names. These measures are calculated using the two corpora for each cluster: the names used by the reviewers and the names used by ChatGPT. Consistency is measured by counting the number of times the most frequently used word appears in each name, while distinctiveness is calculated using the mean Jensen Shannon divergence (see Methods). We see in Fig. 8 that humans and ChatGPT are broadly consistent with each other. The LLM clusters are given the most distinctive names and are also the most consistently interpreted. We see however some interesting differences between the human and machine approaches. ChatGPT is more consistent in its naming between runs, which is unsurprising given it is a single large language model as compared to a set of different human reviewers. This leads to a set of higher interpretability scores when compared to the human scores. However, while the LLM has scored well in both absolute and relative terms using ChatGPT, the random clusters have also performed well in absolute terms. This is likely a consequence of the underlying political skew to the dataset. ChatGPT appears to have better identified this underlying bias in the dataset, by producing lower scores in the distinctiveness measure for the random clusters when compared with the scores given by the human reviewers. An interesting future direction would be to consider datasets with less of an underlying signal.
Figure 8: The interpretability and distinctiveness for each cluster as determined by the names given by human reviewers (circles) and ChatGPT (squares) for the four different clusterings (Doc2Vec, LDA, Random and LLM). The fainter symbols are results by cluster, the more solid symbols with error bars indicate the averages. We can see that the LLM produces the most interpretable and distinctive clusters and that ChatGPT is broadly consistent with the human reviewers, though with significant differences, as discussed in the text.
Discussion
Our work reveals an interesting contrast between reviewer confidence in naming, and the interpretability measure based on naming consistency. When we ask human reviewers to rate their confidence that the name they have provided for a cluster is representative of the cluster as a whole, we see that the LLM-based clusters performed significantly better than the other clustering approaches. However, when we look at consistency between names as provided by the reviewers, we see that the methods are rated more closely, though the LLM clusters again performed the best. This observation is consistent with the findings by Doogan and Buntine, that looking at the results of reviewers completing a labelling task is more informative than asking reviewers about their ability with the task [9]. Focusing on labels also allows us to develop an automated approach based on ChatGPT. Asking ChatGPT to rate its confidence in naming would seem to be of questionable value, but asking ChatGPT to repeatedly provide a name allows us to obtain a measure of interpretability, using the tool’s strength in text comprehension. The top-word consistency measure we have introduced can be seen as a proxy for the inter-coder reliability measure suggested as a means to quantify interpretability [9].
We have chosen to base our distinctiveness measure on names provided, however an alternative is to use the complete cluster information, as was done in our keyword analysis. Complete information better distinguishes randomly created clusters and the separation of clusters in a semantic space, however a measure based on names better reflects the human perception of the clusters. We have chosen this human-centric approach here, however all the documents in a cluster could be combined and a text dissimilarity measure used to quantify the distinctiveness of the clusters [39].
We see that ChatGPT produces results that are consistent with human reviewers, though with some interesting distinctions. ChatGPT found greater variability between the clustering approaches, and appeared to be significantly better at distinguishing the random clusters than the human reviewers. This suggests that ChatGPT is less prone to the biases known to affect humans when naming clusters, such as the temptation to see patterns in noise. However, we should note that the data set we used had a strong political orientation, such that reviewers naming random clusters as ‘political’ is reflective of an underlying signal rather than noise. This signal however is present in all clusters, and ChatGPT appears to have been more consistent in detecting this inherent bias and so overall giving a lower distinctiveness measure to the random clusters than that given by the human reviewers.
The human difficulty of distinguishing a signal from noise is the reason we haven’t implemented a commonly suggested approach for measuring interpretability: topic or word intrusion [10]. The combination of short text and the inherent ambiguity in the clusters due to the absence of a gold standard, make it difficult for a human reviewer to identify an ‘intruder’ whether at the topic or document level. This is particularly the case with short text bios. The information provided by a user may conceivably belong to multiple clusters, with the optimal cluster difficult for a human to determine. For instance, a bio might have a clear political orientation and clear references to family. This ambiguity reflects the multifaceted nature of human identity [30]. However to be of value, the clusters that are found should reflect some deeper divide in the data.
We find the following groupings of bios on Twitter: artistic, occupation, relational, left-leaning, right-leaning, political, general, quotes, emojis, Twitter. This is contrasted with the earlier findings on general classes of identity: occupation, political, kinship, religion, biosocial, ethnicity, corporal, leisure, sexuality, stigma, esteemed, and other [30]. We see some overlap with this earlier classification, with users for instance identifying via their family connections or occupation. We find greater sub-division in the ‘political’ identity, which is unsurprising given the political nature of the dataset. The artistic bios would seem to overlap either with ‘leisure’ or ‘occupation’ so also appear consistent. The ‘general’ cluster was not picked up by the human reviewers, but was found by ChatGPT, and the existence of this class is consistent with the ‘other’ class found earlier. Interestingly, we find distinct classes which appear to be particular to the social media platform. Large numbers of users chose to use emojis to express themselves, making use of this additional avenue for expression (the ‘emojis’ cluster). Others self-identified with reference to the platform itself, appearing to clearly separate their online and offline identities (the ‘Twitter’ cluster). The remaining ‘Quotes’ cluster appears to be an amalgam of identity classes. It had many quotes from the bible, which is consistent with the ‘religion’ class of self-identity, but there were also non-religious quotes used. This cluster, like the emojis cluster, therefore appears to be more a grouping of means of expression rather than particular content (so, a ‘meta’ cluster). Whether people who choose to express themselves using emojis or quotes show other characteristics in common is an interesting direction for further research.
We see that the large language model approach to text clustering has revealed interesting, and human-interpretable, clusters. Based on the immense training sets underpinning such models, more sophisticated patterns have been found, such as the use of quotes or emojis, in addition to more clearly identifiable classes of human self-identity. These clusters also appear to show expected relationships with each other when projected into a two-dimensional space (see Fig. S2), with occupations sitting near artistic bios, but far from political bios. The commonly used approach of LDA appears to have also performed well in producing distinctive and interpretable clusters, however it suffers from a repeatability failure. Clusterings using different random seeds produce quite different partitions of the data, limiting the ability to make general comments about a particular observation (see Fig. S4). This appears to be less of a problem when using the doc2vec method (see Fig. S5), however doc2vec produces clusters that are difficult for a human to interpret (such as grouping bios that use rare words together).
Based on our results, the recommended approach in short-text clustering is to use a large language model to obtain an embedding which can be clustered. For automated analysis of the resulting cluster we have found that ChatGPT has performed as well, or better, than human reviewers. Where ChatGPT performed poorly was in the identification of means of expression rather than content (i.e. meta elements such as the use of quotes or emojis), however on balance the model has still performed better than human reviewers when faced with these clusters. We suggest therefore that LLMs may be used in both the creation and validation of clusters, providing a means to close the well-identified ‘validation’ gap that is common in cluster analysis [33].
However, it is important to acknowledge a fundamental issue with using large language models: they are largely black boxes, and in the case of ChatGPT, maintained by a private company. A risk is therefore that implementations and training sets can change over time, affecting the results. We see this when we perform the cluster interpretation using different ChatGPT implementations (see Fig. S7). However, variability is also an issue with human reviewers (see e.g. the outliers in Fig. 6), and there are additional checks that can be used to validate the LLM clustering measures. A quantification of cluster distinctiveness can be performed using all the clustering data, e.g. via keyword analysis, so providing a valuable metric without needing the automated interpretation step of a second large language model. Our analysis also indicates that the top words are a good guide to the nature of the LLM-created cluster. These can be used to perform a human-based interpretation of a cluster as a check against names created by the second LLM. However, to obtain metrics of interpretability and distinctiveness at scale, we have validated that the second LLM performs well. We hope that this opens up a new approach to analysing short text, allowing new insights to be gained from this increasingly important means of human expression.
3 Methods
3.1 Data
The data used in this paper consisted of 38,639 Twitter user bios from users who used the words ‘trump’ or ‘realDonaldTrump’ between 3rd September and 4th September 2020. The user timelines of these users were collected from Twitter using the Twitter API v2 between September 6th 2020 and September 16th 2020. From these user timelines the username and bio for every user was extracted, however, for this research, only the bio was used.
3.2 Clustering
In this study, clustering models using Python 3 were developed. For data preprocessing, emojis were converted to their CLDR Short Name using the Emoji package [40]. For all models except the Transformer, stopwords were removed and words were lemmatized with NLTK 3.6.5 [41]. LDA and Doc2vec models were built using Gensim 4.1.2 [42]. The Doc2vec method used the parameter values suggested by [12]: 100 dimensions and 75 epochs. The Transformer models employed the all-MiniLM-L6-v2 model [43, 44]. This particular transformer was used as it gives a high performance across different metrics while being fast to run, and has only 384 dimensions when it vectorises text. [44]. A Gaussian mixture model with diagonal covariance was applied to both Doc2vec and Transformer embeddings. Diagonal covariance was chosen as each dimension was orthogonal to each other. The chosen number of clusters and topics (K) was set to 10, a balance between complexity and interpretability.
3.3 Evaluation of model stability
To evaluate the stability of each clustering method, we initially executed each model using 50 distinct random seeds. For every pairwise comparison, we calculated the Adjusted Mutual Information (AMI) between each pair of seeds within the same model. Following this, we computed the overall average AMI and the standard deviation for each model.
3.4 Human Evaluation of Models
Human reviewers (N=39) evaluated clusters/topics generated by LDA, Doc2vec, Transformer, and a randomly created model. Reviewers were recruited from Prolific and were screened to be American and speak English as their first language [45]. From Prolific they were directed to Redcap [46, 47] hosted at the University of Sydney where they saw the data and filled in their responses. For the LDA model, samples included the top 10 high-probability words and 20 bios assigned to each topic. The Doc2vec, Transformer, and random models provided the top 10 frequent words and 20 bios per cluster. Reviewers were shown each cluster from one of the 4 models in a random order, and were asked the following questions:
  • “Create a name using less than 10 words to summarize the top 10 words/emojis and the sample bios of the Sample Cluster. If you believe it is not possible to do this with a cluster, write ‘None’.”
  • “When you named a Sample Cluster, were you confident that the name summarized the whole cluster?” Reviewers can answer this question by selecting one of the following: Not at all Confident/Not Confident/Neutral/Confident/Very Confident.
    For the following questions, reviewers can respond with: Strongly Disagree / Disagree / Neutral / Agree / Strongly Agree
  • “‘The situation when the top words/emojis of a cluster fit together in a natural or reasonable way’ Do you agree that the above statement describes the top words/emojis of this Sample Cluster?”
  • “‘The situation when the Sample bios of a cluster fit together in a natural or reasonable way’ Do you agree that the above statement describes the sample bios of this Sample Cluster?”
  • “‘The Top Words/Emojis provide an accurate summary of the sample bios’ Do you agree that the above statement describes this Sample Cluster?”
See figure S10 and S11 to see how this was presented to the reviewers.
3.5 Ordinal Regression
Survey responses were scaled from 1 (“Strongly Disagree” or “Not at All Confident”) to 5 (“Strongly Agree” or “Very Confident”). The study employed four Bayesian cumulative ordinal regression models, with the dependent variable being the questions the reviewers were asked and the independent variable being the cluster-creating model. All ordinal regression models were created using the BRMS package in R [48]. We used four chains with 5,000 iterations, 2,500 of which were warm-up. This resulted in a total of 10,000 draws from the posterior distribution for each of the four models. The priors for each model were the default from the BRMS package: student’s t-distribution with 3 degrees of freedom, location parameter 0, and scale parameter 2.5. These models generate a latent variable Ŷ [34], with thresholds for each response level. A higher Ŷ indicates a greater likelihood of “Agree” or “Strongly Agree”. Model performance is evaluated by comparing their coefficient distributions against the intercept (random model), set at 0, to assess how much better they perform.
3.6 Keyword Analysis
In keyword analysis, we identify words that occur more frequently in a specific corpus compared to a standard ‘reference corpus,’ as outlined by Baker (2006) [49]. Initially, we count how often each word appears within individual clusters of the corpus. Next, we determine the frequency of these words across the entire corpus. Using these frequency counts, we calculate the Bayes factor for each word in each cluster. Words with a Bayes factor exceeding 10 are classified as keywords. This threshold indicates that a word is significantly more prevalent in a particular cluster compared to its general frequency in the entire corpus [50].
3.7 Metric Calculations
Coherence
Before calculating coherence, all stopwords and punctuation were removed, and every word was lemmatized using NLTK 3.6.5 [41]. Coherence for every cluster was calculated using both the $C_{UMASS}$ and $C_{V}$ methods using the Gensim 4.1.2 package in Python [42]. The top 10 words for each cluster were used in the calculation of coherence.
3.7.1 Distance metrics
For all four models, we utilized the embeddings generated from all-mini-lml, calculating the silhouette score using cosine distance for each individual bio within these models (sklearn [51]). Subsequently, we computed the mean silhouette score for the bios in each respective cluster, providing an aggregated measure of their cohesion. Cluster distance was calculated by finding the mean position of each cluster, then getting the pairwise distance between each cluster within a model. The shortest distance for each cluster was then taken as a measure of how well the bios in that group were clustered. Mean standard deviation for each cluster was calculated similarly to cluster distance, but instead of taking the mean of the cluster, it takes the mean standard deviation across all 384 dimensions.
3.7.2 Correlation between reviewer answers and automated metrics
To test how correlated each automated method was with reviewer responses, for each of the four questions we computed the Spearman rank correlation between all 39 reviewers and all 40 clusters. The Spearman rank correlation, a non-parametric measure, assesses the strength and direction of association between two ranked variables. This gives us a distribution of 39 reviewers and their correlation with the automated metrics. Spearman rank correlation was used since the data is ordinal and a rank-based approach gives the most accurate representation of the data [52].
3.8 Cluster Names
By looking at the names our reviewers gave each cluster, as well as looking at the top 10 words, and reviewing a large number of bios, the authors of this paper created names for each cluster for clarity’s sake. If no meaningful name could be created then the cluster would be designated “U” for “Undecided” and a number, the number serves only as a unique identifier. Clusters within the same model that seemed to warrant the same name (e.g. ‘Left leaning’) were given a 1 or a 2 as a unique identifier.
In our study, a cluster’s names, as provided by our reviewers, were compiled into a corpus. We quantified the frequency of each word within these corpora, subsequently identifying the top 5 most frequently occurring words for every cluster. Additionally, to assess the relative prominence of these words, we converted their counts into percentages based on the number of reviewers (i.e. 39). These percentages were then graphically represented.
We replicated this methodology for the ChatGPT version, with a key modification in data collection. Instead of human reviewers, we employed the ChatGPT API [53], inputting a script that instructed ChatGPT to create a name using up to 5 words for a given cluster. This script included the cluster’s top 10 words and a random sample of Twitter bios for context. In cases where ChatGPT deemed a cluster unnameable or just a random amalgamation, it was instructed to return ‘None’. This procedure was repeated 39 times to match the number of human reviewers and plotted in the same way.
3.9 Distinctiveness and naming measure
To understand the distinctiveness of each cluster, we compile all the names given to a particular cluster into a single corpus, by both ChatGPT and our human reviewers. We apply stemming to all words in these corpora to reduce them to their base forms. Next, we use the Jensen-Shannon divergence to measure the linguistic diversity within each cluster [54]. This is given by $JSD(P\parallel Q)=\frac{1}{2}D_{KL}(P\parallel M)+\frac{1}{2}D_{KL}(Q\parallel M)$. Where $D_{KL}$ is the Kullback-Leibler Divergence. For this, we calculate the JSD between each cluster within a model and identify the smallest distance for each one. We then compute the average and standard deviation of these minimum distances. A higher average Jensen-Shannon divergence suggests that the clusters in our model encapsulate a broader range of topics and exhibit more linguistic variety. Additionally, we created an interpretability metric based on consistency of naming, defined as
[…]